OpenStudy (anonymous):

What are the solutions of 3x2 – 6x – 5 = 0?

OpenStudy (anonymous):

x = 1/3 (3-2 sqrt(6)) x = 1/3 (3+2 sqrt(6)) a = 3 b = -6 x = -5 Use the quadratic formula.

hero (hero):

\[x = \frac{-(2\sqrt{6}-3)}{3}, \frac{2\sqrt{6}+3}{3} \]